### v1

@@ Abstract
Large language models (LLMs) have emerged as pivotal contributors in contemporary natural language processing and are increasingly being applied across a diverse range of industries. However, these large-scale probabilistic statistical models cannot currently ensure the requisite quality in professional content generation. These models often produce “hallucinated” text, compromising their practical utility in professional contexts. To assess the authentic reliability of LLMs in text generation, numerous initiatives have developed benchmark evaluations for hallucination phenomena. Nevertheless, these benchmarks frequently utilize constrained generation techniques due to cost and temporal constraints. These techniques encompass the use of directed hallucination induction and strategies that deliberately alter authentic text to produce hallucinations. These approaches are not congruent with the unrestricted text generation demanded by real-world applications. Furthermore, a well-established Chinese-language dataset dedicated to the evaluation of hallucinations in text generation is presently lacking. Consequently, we have developed an $\mathbb{U}$nconstrained $\mathbb{H}$allucination $\mathbb{G}$eneration Evaluation ($\mathbb{UHG}$Eval) benchmark, designed to compile outputs produced with minimal restrictions by LLMs. Concurrently, we have established a comprehensive benchmark evaluation framework to aid subsequent researchers in undertaking scalable and reproducible experiments. We have also executed extensive experiments, evaluating prominent Chinese language models and the GPT series models to derive professional performance insights regarding hallucination challenges.
Index Terms:
large language models, llms, hallucination, benchmark, unconstrained generation
I Introduction
With the proliferation of extensive textual corpora, the advent of high-performance GPUs, and the refinement of advanced deep learning paradigms, Large language models (LLMs) have exhibited unparalleled proficiency in a multitude of natural language processing (NLP) tasks, including language generation, knowledge application, and intricate reasoning. Concurrently, noteworthy advancements have been realized in the domains of human alignment, engagement with external environments, and the manipulation of tools  [1].
@@ I Introduction
However, LLMs invariably manifest hallucinations [2]. Hallucination is characterized by generated content that is incongruent with user input, the model’s own output context, or factual information. Real-world examples of hallucination from our $\mathbb{UHG}$Eval dataset can be observed in Fig. 1.
@@
Figure 1: Real-world hallucination examples from $\mathbb{UHG}$Eval. Using the IDs, you can locate the corresponding original Chinese news articles within our dataset. Note: MOTIE denotes Ministry of Trade, Industry, and Energy.
@@
Figure 2: The process of creating $\mathbb{UHG}$Eval. Steps 1 to 4 regarding the creation of the benchmark dataset are explained in Section II; Step 5, concerning the evaluation framework, is detailed in Section III.
Owing to reliability concerns, these circumstances markedly hinder the practical deployment of LLMs. Furthermore, in specialized domains like medicine, law, finance, and journalism, hallucination presents a significant impediment to deployment [3, 4]. These fields require stringent standards of content timeliness, accuracy, and logical consistency, attributable to their dynamic and highly specialized characteristics. During the training data collection phase, LLMs may exhibit a deficiency in domain-specific knowledge, yielding outdated content. In the pre-training phase, constraints in model parameters or training methodologies may engender parameter inaccuracies, thwarting the retrieval of accurate content. During the supervised fine-tuning phase, incongruent datasets might yield excessively positive incorrect responses. In the inference phase, the absence of a rollback mechanism can precipitate a cumulative escalation of hallucinations, undermining the logical integrity of responses [5]. For example, erroneous medical guidance, imprecise legal stipulations, and fabricated journalistic narratives substantially restrict the practical utility of LLMs in real-world contexts [3]. The fabricated news content depicted in Fig. 1 offers NO utility to journalists; on the contrary, the verification and rectification of such content exacts a toll on the valuable time of journalists.
Achieving professional-level generation necessitates confronting the significant challenge of devising novel training methodologies and model architectures. However, prior to these developments, it is crucial to formulate a comprehensive, stringent, and demanding benchmark for the assessment of hallucination in language generation [5, 3]. Without such a benchmark, conducting a comparative evaluation of efforts aimed at controlling hallucination would prove to be arduous.
@@
While there have been initiatives to develop benchmarks for hallucination assessment, the majority of these methods employ restricted techniques to produce particular kinds of hallucinated utterances. This approach to generation is at odds with real-world scenarios where hallucinations may arise in unrestricted, spontaneously generated content. For example, HaluEval specifies the type of hallucination in the prompt when generating hallucinated text: “You are trying to answer a question but misunderstand the question context and intention” [6]. Additionally, benchmarks such as HADES annotate hallucinations at a finer granularity by generating token-level hallucinations based on text perturbations [7], but the text perturbation method is still constrained. Ultimately, the majority of benchmarks are centered on the evaluation of hallucinations in English, neglecting the assessment of such phenomena in Chinese. The extensive lexicon of Chinese characters, combined with the complexities introduced by Chinese word segmentation, renders the Chinese hallucination evaluation particularly arduous and deserving of focused scrutiny.
To address the aforementioned challenges, we introduce a novel benchmark for hallucination assessment, as depicted in Fig. 2. The benchmark dataset is comprised of news articles. Selecting texts from this domain is intentional, given that news requires utmost precision in conveying factual information and exhibits minimal tolerance for hallucinations. Constructing an evaluation dataset within this sphere presents a considerable challenge for the majority of LLMs. Concurrently, news articles are of exceptional quality, readily available, and frequently employed as training corpora by a large number of LLMs, guaranteeing impartiality in the evaluation of many LLMs [1]. In light of these factors, we collected a considerable volume of raw news articles, established an efficient, professional-grade hallucination assessment dataset, and formulated an evaluation framework named $\mathbb{UHG}$Eval. It is significant to note that our dataset was produced in an entirely unconstrained fashion. We permit models to compose freely and subsequently sift through the content to identify hallucinations.
@@
Our contributions are as follows: (1) The development of an unconstrained hallucination evaluation dataset. Existing methods for constructing datasets often yield biases towards predefined directions, thereby hindering the full simulation of real-world hallucinations. We have created a hallucination evaluation dataset comprising over 5000 items, generated without intervention, closely mirroring real-world scenarios. (2) The establishment of a unified and diverse evaluation framework. Current benchmark methods for hallucination evaluation often exhibit a singular approach and lack task specificity. We have developed $\mathbb{UHG}$Eval, a unified, flexible, and robust evaluation framework that encompasses generative, discriminative, and selective modalities, along with sentence-level and keyword-level granularity. (3) A comprehensive empirical analysis. We conducted detailed experiments with the proposed benchmark on eight prominent Chinese LLMs and three classic GPT series models to explore the credibility of various LLMs. The aforementioned dataset, evaluation framework, and empirical results collectively constitute the $\mathbb{UHG}$Eval benchmark, which is openly available on Github [^1] https://github.com/IAAR-Shanghai/UHGEval.
II The UHGEval Benchmark Dataset
II-A Data Collection and Pre-processing
@@
To enhance the authenticity of the news continuation dataset, we amassed tens of thousands of historical news articles from leading Chinese news websites, covering the period from January 2015 to January 2017, to serve as the foundation for constructing the dataset. It is worth noting that the decision to eschew the inclusion of more recent news articles (e.g., from 2023) was made to better assess the model’s understanding of existing knowledge and past news events. Indeed, the knowledge embedded within the training data of existing Chinese LLMs typically encompasses information pertaining to significant news between 2015 and 2017  [1].
Considering the different categories of news, such as sports, education, science, and society, the generated hallucinations typically exhibit certain differences. Therefore, when curating the initial news collection for continuation, we endeavored to ensure that the distribution of the collection aligns with the original distribution by randomly sampling from the entire news dataset. Furthermore, we have categorized the collected news examples into four major types: document-intensive, number-intensive, knowledge-intensive, and general news, as shown in Table I. We hypothesize that the likelihood of generating hallucinations varies for different types of news. For example, number-intensive news frequently contains various numerical data, such as years, scores, and values, which may predispose the model to fabricating numbers or introducing minor deviations. Document-intensive news, on the other hand, primarily references official documents, such as factual policy documents, official statements, standard explanations, and legal clauses. In this case, the model may be inclined to fabricate specific policy or document names, or create detailed but fictional policy content. Knowledge-intensive news is characterized by an emphasis on enduring truths and analytical reasoning, which can render the model susceptible to flawed reasoning or the retrieval of incorrect facts. In addition to these three types, we also categorize culturally relevant general news as a separate category for experimental control.
TABLE I: Statistics of Collected News
Type Categories Proportion
DOC Politics, Law, Military, Education 27.52%
NUM Sports, Economy, Market 43.34%
KNO Science, Technology, Healthcare 6.55%
@@
22.59%
  • Note: In the table, DOC denotes document-intensive news; KNO demotes knowledge-intensive news; NUM denotes number-intensive news; GEN denotes general news. The same as below.
@@
Filtering Settings. To ensure the overall quality of the final evaluation dataset, we have implemented the following filters: We consider only the categories listed in Table I, which correspond to the most frequently occurring categories in the original news collection. For news length, we set parameters such that the body length of the selected news falls between 630 and 870 characters, while the beginning text spans between 80 and 120 characters and consists of 2 to 5 sentences. These length parameters reflect the average values in the original news collection and were chosen to avoid overburdening the annotation process at a later stage.
II-B Unconstrained Hallucination Generation
Historically, benchmarks for evaluating hallucination have predominantly relied on a single LLM to produce hallucinated dataset. Notable examples include HaluEval [6] and PHD [8], which exclusively utilize ChatGPT, and FActScore [9] and FACTOR [10], which solely employ InstructGPT [11]. In contrast, our methodology incorporates a suite of five distinct Chinese LLMs to generate hallucinated content. These models include ChatGLM2-6B [12], Baichuan2-13B [13], Qwen-14B [14], InternLM-20B [15], and the Xinyu series model, Xinyu-7B. Xinyu-7B is an augmented large-scale language model derived from the foundational BloomZ-7B [16] through continued pre-training, news-specific fine-tuning, and alignment optimization. Furthermore, Xinyu2-70B is developed based on the open-source LLaMA2-70B [17] framework, incorporating expansions to the Chinese lexicon, ongoing pre-training, and news-specific fine-tuning, thereby endowing it with a robust foundational capability in the news domain. The Xinyu series models are the results of a collaborative research and development effort between the Institute for Advanced Algorithms Research, Shanghai (IAAR, SH), and the State Key Laboratory of Media Convergence Production Technology and Systems of the Xinhua News Agency. Xinyu-7B and Xinyu2-70B will also be utilized in the experiment phase.
Our approach engenders a more heterogeneous generation of hallucinations, mitigating the bias that may arise from the use of a single model and promoting equity within the dataset. This is due to the varying architectures and training corpora inherent to different LLMs. Furthermore, we have adopted an unconstrained generation methodology for the continuation of natural language content. This entails directly inputting the text to be continued into the model without any restrictive prompt instructions, thereby obtaining organic results. For each input example, we concurrently generate five candidate continuations. To maintain consistency across all models, we employ uniform parameter settings, with a temperature coefficient set at 1.0 and max_new_tokens limited to 1024.
II-C Hallucination Ranking
@@
Given the unconstrained nature of our generation paradigm, the task of discerning whether the generated content is indeed hallucinated presents a significant challenge. Upon generating the continuations, a straightforward reliance on human verification is infeasible. An exclusive dependence on human annotation would incur substantial costs and may not be sustainable at scale, whereas a purely machine-based approach, such as utilizing GPT4, could potentially yield less accurate results.
@@
To navigate these complexities, we have adopted a two-stage annotation methodology. This approach begins with an initial phase of hallucination ranking, which is designed to preliminarily sort the generated content based on the likelihood of hallucination. The ranking is then followed by a combination of automatic labeling and human recheck. The integration of hallucination ranking and machine labeling serves a pivotal role in streamlining the subsequent human verification process. This hybrid approach aims to enhance the efficiency and accuracy of human checks, effectively bridging the gap between the scalability of automated processes and the critical discernment of human judgment.
Hallucination ranking is a crucial step in the process of evaluating and selecting the most appropriate continuation from a set of candidate continuations generated by LLMs. The objective of this step is to identify a continuation that not only demonstrates high quality in terms of coherence and readability but also includes an appropriate level of hallucination — misinformation or fabrications that are not supported by the input or real-world knowledge.
To strike this balance, the selection process takes into account two primary dimensions:
@@
Fluency. This refers to the naturalness and readability of the text. A fluent text should read smoothly, be grammatically correct, and make logical sense in the context of the continuation. To assess fluency, a reward model developed by the Institute for Advanced Algorithms Research (IAAR) is employed. This model is trained to evaluate the quality of text and can assign scores to each continuation based on its fluency. By using this model, the top three continuations that exhibit the highest fluency are retained for further consideration.
@@
Likelihood of Hallucination Occurrence. This dimension evaluates the extent to which the continuation may contain hallucinated content. For hallucination occurrence likelihood ranking, we evaluate the lexical correlation between the generated continuation and the reference information. The lower the correlation, the more likely hallucinations are to occur. Despite existing lexical metrics based on n-gram coverage, such as BLEU [18] and ROUGE [19], we believe that these rule-based methods may not effectively discover hallucinated keywords. Therefore, we propose the keyword precision (kwPrec) metric.
@@
This approach initially uses an LLM (here, we use GPT3.5-Turbo) to extract keywords from the continuation and determine whether these keywords have a match in the reference information. The ratio of all matches to the total keywords is then calculated. Since LLMs often extract appropriate keywords more effectively, kwPrec focuses more on factual relevance rather than expressional relevance. Fig. 3 illustrates the tokens segmented by our method compared to those obtained by BLEU-4 and ROUGE-L.
@@
Figure 3: Tokenization results for BLEU-4, ROUGE-L, and kwPrec, using $\text{news}_{\text{id=num\_000432}}$ as an example. The meaning of the above sentence is: Jiangsu is one of the most developed provinces in China for green food production. Note: We ignore tokens that cause overlap.
After implementing this method, we calculate the kwPrec for each of the three candidate continuations, selecting the one with the lowest value as the final candidate. Through the screening in these two stages, we can ensure that, in the worst case scenario, the final candidate continuation ranks third in fluency and third in the likelihood of hallucination occurrence, achieving a balanced level.
By considering both fluency and the likelihood of hallucination, the process aims to filter out continuations that are either too nonsensical or too conservative (lacking any hallucinated content). The ideal candidate continuation would be one that is coherent and engaging but also contains a detectable level of hallucination, which can then be used for further analysis, such as studying the model’s tendencies to hallucinate or for training systems to detect and mitigate such hallucinations.
The final candidate continuations will undergo further annotation to determine the presence and degree of hallucination, which can involve additional automated tools and human judgment. This multi-faceted approach helps ensure that the final selected continuation is both high-quality and relevant for the purposes of the hallucination evaluation benchmark.
II-D Automatic Labeling And Human Recheck

### v2

@@ Abstract
Large language models (LLMs) produce hallucinated text, compromising their practical utility in professional contexts. To assess the reliability of LLMs, numerous initiatives have developed benchmark evaluations for hallucination phenomena. However, they often employ constrained generation techniques to produce the evaluation dataset due to cost and time limitations. For instance, this may involve employing directed hallucination induction or deliberately modifying authentic text to generate hallucinations. These are not congruent with the unrestricted text generation demanded by real-world applications. Furthermore, a well-established Chinese-language dataset dedicated to the evaluation of hallucinations is presently lacking. Consequently, we have developed an $\mathbb{U}$nconstrained $\mathbb{H}$allucination $\mathbb{G}$eneration Evaluation ($\mathbb{UHG}$Eval) benchmark, containing hallucinations generated by LLMs with minimal restrictions. Concurrently, we have established a comprehensive benchmark evaluation framework to aid subsequent researchers in undertaking scalable and reproducible experiments. We have also evaluated prominent Chinese LLMs and the GPT series models to derive insights regarding hallucination.
[…]
$\mathbb{UHG}$Eval: Benchmarking the Hallucination of
Chinese Large Language Models via Unconstrained Generation
Xun Liang*, Shichao Song*, Simin Niu*, Zhiyu Li, Feiyu Xiong, Bo Tang, Yezhaohui Wang, Dawei He, Peng Cheng, Zhonghao Wang, Haiying Deng *School of Information, Renmin University of China, Beijing, China Institute for Advanced Algorithms Research, Shanghai, China State Key Laboratory of Media Convergence Production Technology and Systems, Beijing, China
footnotetext: * These authors contribute equally
Γ Corresponding author: lizy@iaar.ac.cn
1 Introduction
@@
Large language models (LLMs) have unparalleled proficiency in language generation, knowledge application, and intricate reasoning Zhao et al. (2023). However, these models invariably manifest hallucination Rawte et al. (2023), as they often generate content that is incongruent with user input, the model’s output context, or factual information. Real-world hallucination examples from our $\mathbb{UHG}$Eval dataset can be observed in Fig. 1.
@@
Figure 1: Hallucinations from $\mathbb{UHG}$Eval. Using the IDs, you can locate the original news articles. Note: MOTIE denotes Ministry of Trade, Industry, and Energy. (In Chinese: Fig. 10)
@@
Figure 2: The process of creating $\mathbb{UHG}$Eval. Steps 1 to 4 regarding the creation of the benchmark dataset are explained in Section 2; Step 5, concerning the evaluation framework, is detailed in Section 3. (In Chinese: Fig. 11)
The fabricated news content depicted in Fig. 1 offers NO utility to journalists; on the contrary, the verification and rectification of such content exacts a toll on the valuable time of journalists. To this concern, it is crucial to first formulate a comprehensive, stringent, and demanding benchmark for the assessment of hallucination in language generation Zhang et al. (2023); Wang et al. (2023b).
@@
While there have been a bunch of efforts to develop benchmarks for hallucination assessment, they always employ restricted techniques to produce particular kinds of hallucinated utterances. This approach is at odds with real-world scenarios where hallucinations arise in unrestricted, spontaneously generated content. For example, HaluEval specifies the type of hallucination in the prompt when generating hallucinated text: “You are trying to answer a question but misunderstand the question context and intention” Li et al. (2023). Additionally, benchmarks such as HADES annotate hallucinations at a finer granularity by generating token-level hallucinations based on text perturbations Liu et al. (2022), but the text perturbation method is still constrained. Besides, many benchmarks are centered on the evaluation of hallucinations in English, neglecting the assessment of such phenomena in Chinese. The extensive lexicon of Chinese characters, combined with the complexities introduced by Chinese word segmentation, renders the Chinese hallucination evaluation particularly arduous and deserving of focused scrutiny.
To address the aforementioned challenges, we introduce a novel benchmark for hallucination assessment, as depicted in Fig. 2. The benchmark dataset is composed of raw Chinese news articles and continuations of those articles freely generated by LLMs but annotated with hallucinations.
Furthermore, selecting texts from the news domain is intentional, given that news requires utmost precision in conveying factual information and exhibits minimal tolerance for hallucinations, presenting a considerable challenge for the majority of LLMs. Moreover, news data encompasses a wide range of topics, including medicine, technology, finance, sports, etc., incorporating features found in texts from other domains. Lastly, news articles are readily available and frequently employed as training corpora by a large number of LLMs, guaranteeing impartiality in the evaluation of many LLMs Zhao et al. (2023).
@@
Our contributions: (1) The development of an unconstrained hallucination evaluation dataset, comprising over 5000 items. Existing methods for constructing datasets often yield biases towards predefined directions, thereby hindering the full simulation of real-world hallucinations. (2) The establishment of a unified and diverse evaluation framework, $\mathbb{UHG}$Eval, that encompasses discriminative, selective, and generative evaluations. Current benchmark methods for hallucination evaluation often exhibit a singular approach and lack task specificity. (3) A comprehensive empirical analysis. We evaluated eight prominent Chinese LLMs and three classic GPT series models to explore the credibility of various LLMs. The aforementioned dataset, evaluation framework, and empirical results collectively constitute the $\mathbb{UHG}$Eval benchmark, which is openly available on GitHub111 Main: https://github.com/IAAR-Shanghai/UHGEval
Dataset creation: https://github.com/IAAR-Shanghai/UHGEval-dataset
2 The UHGEval Dataset
2.1 Data Collection and Pre-processing
@@
We amassed tens of thousands of historical news articles from leading Chinese news websites, covering the period from January 2015 to January 2017, to serve as the foundation for constructing the dataset. It is worth noting that the decision to eschew more recent news articles (e.g., from 2024) was made to better assess the model’s understanding of existing knowledge. Indeed, the knowledge embedded within the training data of existing Chinese LLMs typically encompasses information about significant news between 2015 and 2017 Zhao et al. (2023).
The collected news spans various topics, such as sports, education, science, society, finance, and more. This diversity underscores the advantage of choosing news texts for our dataset, as it enables the incorporation of a wide array of text genres. We hypothesize that the occurrence of hallucinations will vary as LLMs generate news across different categories. As a result, we have classified these diverse categories into four main types: document-intensive, number-intensive, knowledge-intensive, and general news, with details provided in Table 1.
Type Share Categories
DOC
27.52%
Politics, Law, Military, Education
NUM
43.34%
Sports, Economy, Market
KNO
6.55%
Science, Technology, Healthcare
@@
Table 1: Statistics of collected news. DOC, NUM, KNO, and GEN denote document-intensive, number-intensive, knowledge-intensive, and general news, respectively.
@@
2.2 Unconstrained Hallucination Generation
Unlike directed hallucination generation Li et al. (2023) or perturbation-based generation Liu et al. (2022), we have adopted an unconstrained generation methodology for the continuation of natural language content, though it poses difficulties for subsequent annotations. This generation’s fashion entails directly inputting the text to be continued into the model without any restrictive prompt instructions, thereby obtaining organic results.
Furthermore, current benchmarks for evaluating hallucination have predominantly relied on a single LLM to produce a hallucinated dataset. Notable examples include HaluEval Li et al. (2023) and PHD Yang et al. (2023b), which exclusively utilize ChatGPT, and FActScore Min et al. (2023) and FACTOR Muhlgay et al. (2023), which solely employ InstructGPT Ouyang et al. (2022). In contrast, our methodology incorporates a suite of five distinct Chinese LLMs to generate hallucinated content. These models include ChatGLM2-6B Du et al. (2022), Baichuan2-13B Yang et al. (2023a), Qwen-14B Bai et al. (2023), InternLM-20B InternLM (2023), and Xinyu-7B. For additional information about the Xinyu series models, please refer to the Appendix B.1.
For each input news article, we concurrently generate five candidate continuations using 5 different LLMs without constraint. Overall, our approach engenders a more unconstrained and heterogeneous generation of hallucinations, mitigating the bias that may arise from the use of a single model or constrained prompting.
2.3 Hallucination Ranking
@@
Given the unconstrained nature of our paradigm, the task of discerning whether the generated content is indeed hallucinated presents a significant challenge. Upon generating the continuations, an exclusive dependence on human annotation would incur substantial costs, whereas a purely machine-based approach, such as utilizing GPT4, could potentially yield less accurate results.
@@
To navigate these complexities, we have adopted a two-stage annotation. This approach begins with an initial stage of hallucination ranking (Section 2.3), designed to sort the generated content based on the likelihood of hallucination. The ranking is then followed by the second stage of automatic labeling and human rechecking (Section 2.4).
Algorithm 1 Hallucination ranking
$candidate:list[str]$
$final:str$
$candidate.sort(descend,by=fluency)$
$picked\leftarrow candidate[:3]$triangleright More fluency
$picked.sort(descend,by=likelihood)$
$final\leftarrow picked[0]$triangleright More Hallucination
Hallucination ranking is a crucial step in selecting the most appropriate continuation from a set of candidate continuations generated by LLMs. We employ a ranking process detailed in Algorithm 1. This process relies on two critical metrics: $fluency$, ensuring that the continuation does not become too nonsensical, and $likelihood$, which stands for the likelihood of hallucination occurrence, ensuring that the continuation includes a detectable level of hallucination. By employing such a ranking, it is guaranteed that, in the worst-case scenario, the $final$ candidate ranks at least third in fluency and third in the likelihood of hallucination occurrence, achieving a balanced level. The two metrics are computed as follows.
Fluency
@@
This refers to the coherence and readability of the text. A fluent text should read smoothly, be grammatically correct, and make logical sense in the context of the continuation. To assess fluency, a reward model developed by the Institute for Advanced Algorithms Research (IAAR) is employed, trained to score text quality based on fluency.
Likelihood of Hallucination Occurrence
@@
This dimension evaluates the extent to which the continuation may contain hallucinated content. To estimate the probability, we evaluate the lexical correlation between the generated continuation and the reference information. The lower the correlation, the more likely hallucinations are to occur. Despite existing metrics, such as BLEU Papineni et al. (2002) and ROUGE Lin (2004), we believe that these rule-based methods may not effectively discover hallucinations. Therefore, we propose the keyword precision (kwPrec) metric.
@@
This metric uses an LLM (e.g., GPT3.5-Turbo) to extract keywords from the continuation and determine whether these keywords have a match in the reference information. The ratio of all matches to the total keywords is then calculated. Since LLMs often extract appropriate keywords more effectively, kwPrec focuses more on factual relevance rather than expressional relevance. Fig. 3 illustrates the tokens segmented by our method compared to those obtained by BLEU-4 and ROUGE-L.
@@
Figure 3: Tokenization results for BLEU-4, ROUGE-L, and kwPrec, using $\text{news}_{\text{id=num\_000432}}$ as an example. The meaning of the above sentence: Jiangsu is one of the most developed provinces in China for green food production.
2.4 Automatic Labeling and Human Rechecking